OpenStudy (calculusxy):

A bug is moving along the line 12x + 5y = 60 with constant speed 5 units per second. It crosses the x-axis when t = 0 seconds and crosses the y-axis later. When? What vector represents the movement of the bug each second?

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

@SolomonZelman @mathstudent55

OpenStudy (agent0smith):

First draw a graph cos that'll help. Find the slope, draw a triangle, make the hypotenuse 5. Then you can use the slope to write expressions for the other sides.

OpenStudy (calculusxy):

The slope is -12/5. Why would the hypotenuse be 5?

OpenStudy (calculusxy):

Wouldn't it be 13?

OpenStudy (calculusxy):

Would it be correct that the bug crosses the y-axis when t=1?

OpenStudy (agent0smith):



OpenStudy (calculusxy):

y = 12 - (12/5)x Solving for the x-axis 0 = 12 - (12/5)x -12 = -(12/5)x 5 = x When t = 0, the bug crosses (5,0) Solving for the y-axis y = 12 - (12/5)(0) y = 12 (5,0) (0,12) x = 5 - 5t y = 12t When x=0 then it crosses the y-axis 0 = 5 - 5t -5 = -5t 1 = t When t=1, the bug crosses the y-axis?

OpenStudy (agent0smith):

That doesn't look right. It travelled 13 units in 1 second. See my above post to work out speed.

OpenStudy (agent0smith):